OpenStudy (anonymous):

11. The formula for converting degrees Fahrenheit to degrees Celsius is F = 9/5C + 32. At what Celsius temperature is F = C?

OpenStudy (kunal):

-40................... put F=C and solve the equation

OpenStudy (anonymous):

this is how approached the problem 5*F= (9/5C + 32)* 5 **I multiplied 5 to cancel out the fraction ____________________ 5 5 F=9C+32 __________ 5 Now I am conflicted

OpenStudy (anonymous):

17. Which of the following are true? (i) 3 dam=3*10^3 cm (ii) 22mm=22*10^6 km (iii) 7 dm=7 10^-2 dam A) (i) and (ii) only B) (i) and (iii) only C) (ii) and (iii) only D) (i), (ii), and (iii)

OpenStudy (callisto):

Since F=C F = 9/5C + 32 C = 9/5C + 32 -32 = 4/5C C = -32 /(4/5) = -40

OpenStudy (callisto):

1 dam = 1 000 cm => 3 dam = 3 000 cm = 3x 10^3 cm 1km = 1000m = 100 000 cm = 1 000 000mm 1dm =10cm Can you try?

OpenStudy (anonymous):

yes. thank you for your help.